What must the surface area of a piston be,if a pressure of 300kPa is used to give a desired upward force of 2000N?


1
Expert's answer
2021-06-03T09:15:11-0400

By the definition of the pressure, we have:


"P=\dfrac{F}{A},""A=\dfrac{F}{P}=\dfrac{2000\ N}{3\cdot10^5\ \dfrac{N}{m^2}}=6.67\cdot10^{-3}\ m^2=66.7\ cm^2."